OpenStudy (anonymous):

A constant force F = 1i - 7j +7k moves an object along a straight line from point (2,7,-5) to point (5,3,-10). Find the work done if the distance is measured in meters and the magnitude of the force is measured in newtons.

OpenStudy (anonymous):

F= 1i - 7j + 7k W=F.S W=(1i - 7j + 7k)(3i -4j -5k) W= 3 + 28 -35

OpenStudy (anonymous):

W= -4 Nm

OpenStudy (anonymous):

oh that's all...don't have to use the cross product?

OpenStudy (anonymous):

No...Work is Scalar No Cross Product ....Dot Product is what we need here

OpenStudy (anonymous):

well then duh..that was simple

OpenStudy (anonymous):

Yesh!